Determine the angle between the following two vectors.

|𝑔| = (1,2,−3)


|h| = (0,−3,4)


1
Expert's answer
2021-04-29T18:06:56-0400

Let "\theta" be the angle between the vectors "\vec{g}" and "\vec{h}" .

Using the transformed dot product equation,


"\theta = \cos^{-1} \left( \frac{ \vec{g}\cdot \vec{h}}{|\vec{g}| |\vec{h}| } \right)"

where

"\vec{g}\cdot \vec{h} = (1)(0) + (2)(-3)+(-3)(4)=-18" (the dot product of "\vec{g}" and "\vec{h}" )

and

"|\vec{g}| = \sqrt{(1)^2+(2)^2+(-3)^2} = \sqrt{14}" (magnitude of vector "\vec{g}" )

"|\vec{h}| = \sqrt{(0)^2+(-3)^2+(4)^2} = \sqrt{25}=5" (magnitude of vector "\vec{h}" )


Therefore,


"\theta = \cos^{-1} \left( \frac{-18}{\sqrt{14} (5)} \right)"

"\theta = \cos^{-1} \left( -0.9621 \right)"

"\theta = 164.18^{0}"